11. A college student is trying to decide whether to buy or lease a motorcycle for getting to and from school. The cost of the motorcycle is $13,000, including tax.

a. Determine the monthly payment on a loan for the bike with a financing term of 36 months, an APR of 5.9%, compounded monthly.

b. Determine the monthly payment on a lease for 24 months, an APR of 4.1%, compounded monthly. The deprecation is 15% per year.

c. Based solely on the monthly payment, which option is the most affordable for this student.
Answer #1

Monthly payment = \small \frac{Principal * r * (1+r)^{n}}{(1+r)^{n}-1}

r = 0.059 / 12 = 0.00491666666

n = 36 months

= \small \frac{13000 * 0.00491666666 * (1+0.00491666666)^{36}}{(1+0.00491666666)^{36}-1}

= 394.90

Monthly payment = \small \frac{Principal * r * (1+r)^{n}}{(1+r)^{n}-1}

r = 0.041 / 12 = 0.00341666666

n = 24 months

= \small \frac{13000 * 0.00341666666 * (1+0.00341666666)^{24}}{(1+0.00341666666)^{24}-1}

= 1.08530411145

= 565.10

Option A is better in terms of affordability.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.